OpenStudy (anonymous):

Fill in the blank to make the expression a perfect square. w^2-12w+?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Again, split the middle term in half\[\large\rm w^2\color{orangered}{-12w}\quad=w^2\color{orangered}{-6w-6w}\] And then square this value (multiply it by itself) to get your answer. (-6)^2 = ?

OpenStudy (anonymous):

36.

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

So if I had x^2+10x+? It would be 25^2?

OpenStudy (anonymous):

5* not 25 lol

zepdrix (zepdrix):

Ah good good, there we go. 5^2 which is 25, not 25^2 hehe ^^

OpenStudy (anonymous):

What if I had x^2-2x+? It would be 1 right?

zepdrix (zepdrix):

Good good. (-1)*(-1) = 1